OpenStudy (anonymous):

WILL FAN AND MEDAL!!! Simplify sine theta over square root of the quantity 1 minus sine squared theta.

OpenStudy (anonymous):

@ganeshie8 @bibby

OpenStudy (anonymous):

@texaschic101 @dannyrod2000

OpenStudy (anonymous):

@kropot72

OpenStudy (driftracer305):

\[(\sin \theta )\div (1 - \sin^2 \theta )\] ??

OpenStudy (driftracer305):

ohh k sq root....got it...wait

OpenStudy (driftracer305):

so sin theta / sqrt. cos^2 theta............ then \[(\sin \theta) \div (\cos \theta) = \tan \]

OpenStudy (driftracer305):

@cambam1323

OpenStudy (driftracer305):

@cambam1323 ...........u got it?

OpenStudy (anonymous):

no. these are the answer choices.... tan Θ tan2 Θ 1 −1

OpenStudy (anonymous):

sorry, its tan theta tan^2 theta 1 -1

OpenStudy (driftracer305):

yea.....i meant tan theta

OpenStudy (driftracer305):

so it's A

OpenStudy (anonymous):

okay thank you!

OpenStudy (driftracer305):

mah pleasure